OpenStudy (anonymous):

The surface area of a cone with radius r units and slant height s units is shown below: Surface area = 3.14(rs + r2) Part A: If r = 3 units and s = 5 units, write an expression that can be used to calculate the surface area of the cone. (4 points) Part B: What is the surface area of the cone? Show your work. (6 points)

OpenStudy (anonymous):

@Holly00d1248 @UsukiDoll @dtan5457

OpenStudy (anonymous):

@Nnesha @robtobey

Nnesha (nnesha):

just substitute r and s values into the formula then solve

OpenStudy (anonymous):

? @Nnesha

Nnesha (nnesha):

they gave u the formula just replace r by 3 and s by 5 then solve

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

15?

Nnesha (nnesha):

nope

Nnesha (nnesha):

if you meant r times s =15 then yes

Nnesha (nnesha):

\[\huge\rm SA = 3.14(rs+r^2)\] replace r by 3 and s by 5

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@izthewiz7

Nnesha (nnesha):

what grade are you in ?? r u flvs student ?

Nnesha (nnesha):

connection or whatever the name of school ? :)

Nnesha (nnesha):

@jdoe0001 can you help this user please ? i gtg

OpenStudy (jdoe0001):

k

OpenStudy (anonymous):

@jdoe0001 The surface area of a cone with radius r units and slant height s units is shown below: Surface area = 3.14(rs + r2) Part A: If r = 3 units and s = 5 units, write an expression that can be used to calculate the surface area of the cone. (4 points) Part B: What is the surface area of the cone? Show your work. (6 points) @jdoe0001

OpenStudy (jdoe0001):

\(\large { \textit{surface area of a cone}=\pi(rs+r^2)\quad \\ \quad \\ \begin{cases} \pi \to &{\color{blue}{ 3.14}}\\ r\to radius\to &{\color{purple}{ 3}}\\ s\to \textit{slanted height}\to &{\color{brown}{ 5}} \end{cases} \\ \quad \\ {\color{blue}{ 3.14}}({\color{purple}{ r}}{\color{brown}{ s}}+{\color{purple}{ r}}^2)\implies ? }\)

OpenStudy (anonymous):

? @jdoe0001

OpenStudy (jdoe0001):

well... tis just a matter of plugging in the values

OpenStudy (jdoe0001):

sure you've done that already, many times by now